OpenStudy (anonymous):

A sample of gold (Au) has a mass of 35.12 g. a. Calculate the number of moles of gold (Au) in the sample and record in Table 1. Show your work. b. Calculate the number of atoms of gold (Au) in the sample and record in Table 1. Show your work.

OpenStudy (abb0t):

To get the moles, you need to divide by the molar mass. To get atoms, you would multiply by avogadro's constant: \(\sf \color{red}{6.022 \times 10^{23} \frac{atoms}{mol}}\) You should see that units cancel, leaving you with desired units. For instance: \(\sf \color{blue}{a \times \frac{b}{a}} \Rightarrow \cancel{a} \times \large \frac{b}{\cancel{a}} \) see how the two a's cancel. it is the same when you have grams.

OpenStudy (anonymous):

Thank you : ) I am suppose to be doing a portfolio with another problem like that, and I have to fill out a table and I missed so much of this lesson being sick so I am SO confuse! :/ Thanks again!